OpenStudy (anonymous):

MEDAL!An object's motion is described by the equation d = 4 sin(pi t ) . The displacement, d , is measured in meters. The time, t , is measured in seconds. Answer the following questions: (a) What is the object's position at t = 0 ? (b) What is the object's maximum displacement from its resting position? (c) How much time is required for one oscillation? (d) What is the frequency?

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

this is the last one on my study guide :)

OpenStudy (ikram002p):

okeyy lets do it !

OpenStudy (anonymous):

:))))

OpenStudy (ikram002p):

(a) What is the object's position at t = 0 ? so u wanna find d at 0 d = 4 sin(pi t ) find d(0) do u know how to do that ?

OpenStudy (anonymous):

A) is the objects position 0

OpenStudy (ikram002p):

yess ur correct !

OpenStudy (ikram002p):

nw do u know how to find maximum displacement ?

OpenStudy (anonymous):

isn't it the total distance the object traveled for this

OpenStudy (ikram002p):

yess it is , so do u know how to find it ?

OpenStudy (anonymous):

not here I don't, so I could use some help

OpenStudy (ikram002p):

ok do u know how to find d' ?

OpenStudy (anonymous):

only if I know t

OpenStudy (ikram002p):

diffrentiate d ?

OpenStudy (ikram002p):

no u dnt need to know t , u wanna find t

OpenStudy (anonymous):

then I don't know how to find d

OpenStudy (ikram002p):

maximum displacment acures when d'=0

OpenStudy (ikram002p):

ok ok ill show u

OpenStudy (anonymous):

ok

OpenStudy (ikram002p):

d = 4 sin(pi t ) d'=4 cos (pi t) do u know this ?

OpenStudy (anonymous):

I think. do i do 0=4cos(pi t)

OpenStudy (ikram002p):

yess ur right ! and find t :)

OpenStudy (anonymous):

t=1/2? or 3pi/2

OpenStudy (ikram002p):

yess ur correct ! so to chose which one is correct we see which of them gives the hight distance in d = 4 sin(pi t )

OpenStudy (anonymous):

so 1/2 is the maximum displacement

OpenStudy (ikram002p):

yess ur correct cuz 3pi/2 gives negative answer

OpenStudy (ikram002p):

but not 1/2 is the displacment 1/2 is the time at the max displacment d(1/2) is the max d(1/2)=4 sin (pi /2) =4 ( so 4 is the max displacment)

OpenStudy (anonymous):

oh ok

OpenStudy (ikram002p):